OpenStudy (anonymous):

Sheila took the temperature in her bedroom at 9:00 p.m. before she went to bed. Overnight, the temperature dropped 9°F but then increased 11°F by morning. During the day, the temperature raised another 8°F. Which integer represents the total change in temperature? -28°F -12°F +10°F +28°F

OpenStudy (ja1):

Ok so let's do a bit of addition and subtraction k: so first: -9 +11 +8 so -9 + 11 = ?

OpenStudy (ja1):

\[-9 + 11 =\]

OpenStudy (ja1):

That is your first question.

OpenStudy (anonymous):

well 9 + 11 = 20

OpenStudy (ja1):

Yes but remember the 9 is negative so if we reorder it like so 11 - 9 =

OpenStudy (anonymous):

oh ok then it would be 2

OpenStudy (ja1):

Yes because -9 + 11 = 2 :D so now do this: \[2 + 8 = ?\]

OpenStudy (anonymous):

it would be 10

OpenStudy (ja1):

YES so it is 10F :D

OpenStudy (anonymous):

ohh ok thanx you

OpenStudy (ja1):

No prob, I see you got the medals down pat :D fast learner

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

can u hlp me with one last question?

OpenStudy (ja1):

Yes just post it here for the sake of time :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A truck driver drives 10 miles west from company headquarters when he realizes he forgot to make one of the stops on his route. He drives 7 miles east on the same road to go back and make the stop. Which expression can be used to find how far the truck driver is from company headquarters? +10 + -7 -7 + -10 +10 + +7 -7 - +10

OpenStudy (ja1):

Ok so he went out 10 miles ut hewent back seven so we need to subtract 7 from ten, which equation would thta be?

OpenStudy (anonymous):

it would be 3

OpenStudy (ja1):

Um not quite, 13 would be 10 + 7 you need 10 - 7

OpenStudy (anonymous):

oh ok

OpenStudy (ja1):

Remember what I said about changing out the signs for words(minues) (negative)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok I its 17 right

OpenStudy (ja1):

Oh you mean the product, I am talking about which equation lol

OpenStudy (anonymous):

lol

OpenStudy (ja1):

But you were right it was 3 so which equation is that?

OpenStudy (anonymous):

its +10 + +7 right

OpenStudy (ja1):

no that would equal 17, it has to equal 3 like you said before

OpenStudy (anonymous):

oh sorry I forgot about that

OpenStudy (ja1):

xD its fine :)

OpenStudy (anonymous):

lol its the last one

OpenStudy (anonymous):

:)

OpenStudy (ja1):

close but its actually the first one, you just mixed up the numbers :)

OpenStudy (anonymous):

oh ok thank you a lot you helped me out big time lol :)

OpenStudy (ja1):

No prob thtas what this site is all about, if you ever need help adjusting just PM me :)

OpenStudy (anonymous):

ok :) thanx